# Flock: A Knowledge Graph Foundation Model via Learning on Random Walks

### v1

@@ C.2 Invariance
Similarly, let $\eta=v_{0}\xrightarrow{r_{1}}\dots\xrightarrow{r_{l}}v_{l}$ be a walk of length $l$ in $G$. The view of $\eta$ with $\mu$ is defined as:
$\mu\left(v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}\right)=\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\dots\xrightarrow{\phi(r_{l})}\pi(v_{l})$
@@
Following the definition of invariance in probability, provided in Section 3, we extend all the definitions above to the stochastic case, replacing equality (==) with equality in distribution ($\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}$).
We can now prove the main propositions stated in Section 4.2. Let’s begin with the more general:
@@
We can use the conclusion from Proposition 4.2 to prove the probabilistic invariance of the architecture proposed in Section 4. To be able to apply it, we first need to verify the invariance of all used components, which we formalize in the following lemmas.
Lemma C.6.
The choice of the first step $v_{0}\xrightarrow{r_{1}}v_{1}$ of the uniform random walk algorithm described in Section B.1 is invariant.
@@
Let $G=(V,E,R)$ be a graph and let $H\simeq G$ be an isomorphic graph, with the isomorphism $\mu=(\pi,\phi)$ taking $G$ to $H$. Consider a link prediction query $q=(h,r,?)$ over $G$, and its identification $q^{\prime}=\mu(q)=(\pi(h),\phi(r),?)$. The goal is to show that when using $\eta$ described in Section B.1 for $(G,q)$ and $(H,q^{\prime})$, the first steps:
@@
Let $G=(V,E,R)$ be a knowledge graph, and let $l$ be the length of random walks. Let $H$ be a KG isomorphic to $G$ via the isomorphism $\mu=(\pi,\phi)$. We aim to show that:
$\mu(\eta(G,l))=\pi(V_{0})\xrightarrow{\phi(R_{1})}\pi(V_{1})\xrightarrow{\phi(R_{2})}\dots\xrightarrow{\phi(R_{l})}\pi(V_{l})\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}U_{0}\xrightarrow{S_{1}}U_{1}\xrightarrow{S_{2}}\dots\xrightarrow{S_{l}}U_{l}=\eta(H,l)$
Let $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}\in\mathcal{W}(G)$ be a walk of length $l$ over $G$. It suffices to show that the probability of sampling $\bar{\eta}$ from $G$ is identical to the probability of sampling $\mu(\bar{\eta})$ from $H$:
${\mathbb{P}}(\eta(G,l)=\bar{\eta})={\mathbb{P}}(\eta(H,l)=\mu(\bar{\eta}))$
To see this, let us expand the definitions of ${\mathbb{P}}(\eta(G,l)=\bar{\eta})$:
${\mathbb{P}}(\eta(G,l)=\bar{\eta})=$ ${\mathbb{P}}(V_{0}=v_{0})$ (15)
@@
$\cdot\prod^{l-2}_{i=0}{\mathbb{P}}(V_{i+2}=v_{i+2}\mid V_{i+1}=v_{i+1},V_{i}=v_{i})$
$\cdot\prod^{l-1}_{i=0}{\mathbb{P}}(R_{i+1}=r_{i+1}\mid V_{i+1}=v_{i+1},V_{i}=v_{i})$
and $P(\eta(H,l)=\mu(\bar{\eta}))$:
${\mathbb{P}}(\eta(H,l)=\mu(\bar{\eta}))=$ ${\mathbb{P}}(U_{0}=\pi(v_{0}))$ (16)
@@
$\cdot\prod^{l-2}_{i=0}{\mathbb{P}}(U_{i+2}=\pi(v_{i+2})\mid U_{i+1}=\pi(v_{i+1}),U_{i}=\pi(v_{i}))$
$\cdot\prod^{l-1}_{i=0}{\mathbb{P}}(S_{i+1}=\phi(r_{i+1})\mid U_{i+1}=\pi(v_{i+1}),U_{i}=\pi(v_{i}))$
@@
for all $i\in\{0,1,\dots,l-2\},j\in\{1,\dots,l-1\}$. Moreover, by the assumption that the first step $V_{0}\xrightarrow{R_{1}}V_{1}$ is invariant, we have:
@@
Substituting these equalities into (18) and multiplying both sides by the equalities from (17) for all choices of $i\in\{0,1,\dots,l-2\},j\in\{1,\dots,l-1\}$, we get precisely the equality of the right sides of equations (15) and (16). Hence,
${\mathbb{P}}(\eta(G,l)=\bar{\eta})={\mathbb{P}}(\eta(H,l)=\mu(\bar{\eta}))$
and we can conclude that $\mu(\eta(G,l))\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\eta(H,l)$, and the algorithm $\eta$ is invariant in probability. ∎
Corollary C.8.
The random walk algorithm presented in Section B.1 is invariant in probability.
Lemma C.9.
The recording protocol $w$, as described in Section B.2, is invariant, provided that the embedding functions $\mathbf{v}$ and $\mathbf{r}$ are invariant.
@@
Let $G=(V,E,R)$ and $H=(V^{\prime},E^{\prime},R^{\prime})$ be isomorphic knowledge graphs with the isomorphism $\mu=(\pi,\phi)$ taking $G$ to $H$. Let $q=(h_{q},r_{q},?)$ be a link prediction query over $G$, and $\mu(q)=(\pi(h_{q}),\phi(r_{q}),?)$ be its identification in $H$. Let $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{l}}v_{l}\in\mathcal{W}(G)$ be a walk over $G$, and $\bar{\eta}^{\prime}=\mu(\bar{\eta})=\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\dots\xrightarrow{\phi(r_{l})}\pi(v_{l})$ be the analogous walk over $H$. To prove that the recording protocol $w$ outlined in Section B.2 is invariant, it suffices to show that the encoding of each step:
@@
Lemma C.10.
The consensus protocol $c$, as described in Section B.4, is invariant.
@@
Flock with components as described in Section 4 is invariant in probability.
@@
The result follows naturally from aggregating the results of Corollary C.8 and Lemmas C.9 and C.10, followed by applying Proposition 4.2. ∎
@@ Appendix D Computational Complexity
Recall that $I$ is the iterations in each forward pass of Flock; $n$ is the base walk count; $\ell$ is the walk length; $L$ is the number of linear sequence-model layers (such as GRU); and $d$ is the hidden dimension for the sequence processor. Note that in practice, we perform $P$ forward passes and ensemble their outputs to reduce variance. For a single pass (P=1P{=}1), walk sampling and recording cost $O(n\ell)$, while the sequence processor with $L$ layers of hidden dimension $d$ costs $O(n\ell Ld^{2})$. The consensus protocol costs $O(n\ell d)$. In total, the time complexity is $O\big(PIn\ell Ld^{2}\big)$, which scales linearly with the number of (base) walks $n$, the length of walks $\ell$, and the number of ensembled predictions $P$. We empirically verified this in Appendix E.
@@ Appendix E Scalability analysis
Table 6: Inference scalability on a single NVIDIA RTX A6000 (48 GB); batch size =8=8. Left columns specify base walks $n$ and ensembled passes $P$. Dashes indicate not applicable.
@@
To investigate the scalability of the proposed method Flock, we report the training and inference time per batch and peak GPU memory for Ultra, Trix, and variants of Flock on a single RTX A6000 (48 GB) in Tables 5 and 6.
@@
During training, we fix Flock to $n=16$ base walks and with an ensemble size of $P=1$, which yields higher cost than Ultra/Trix but remains feasible on a single GPU. In addition, unlike Ultra/Trix, Flock does not rely on GNN message passing where highly optimized fused sparse kernels (e.g., RSPMM kernel developed in Zhu et al. (2021)) accelerate computation; instead, Flock’s runtime is dominated by walk sampling and sequence encoding, making time per batch the main bottleneck. As a result, pretraining typically takes about three days. One avenue for future work is to develop similarly highly optimized kernels for random-walk sampling to speed up the process.
@@
Additionally, we report the inference results in Table 6, where we vary the number of walks $n$ and ensembled passes $P$. We observe near-linear growth of latency and VRAM with $n$, reflecting the dominant costs of walk sampling and sequence processing. Note that during inference, ensembled predictions are parallelizable, meaning that with sufficient GPU memory, these $P$ stochastic passes can be executed concurrently, so the latency grows sublinearly in $P$, while peak VRAM scales roughly linearly with $P$. In practice, reducing $n$ (walks) or $P$ (ensembled passes) lowers both memory and latency, while larger $n/P$ settings trade extra cost for better coverage and stability on harder KGs.
@@ Appendix F Noise injection over existing KGFMs
Setup.
To answer Q5, we apply noise injection over the best performing KGFMs baselines Trix. Specifically, in each forward pass, we subtract an element-wise noise sampled from a uniform distribution $\epsilon\sim{\mathcal{U}}[0,0.5]$ to all relation and entity embeddings after the initialization stage. Note that the addition of noise will technically break deterministic node-relation equivariance, but the resulting model (Trix + noise) still respects probabilistic node-relation equivariance. We then pretrain Trix using the same experimental setup shown in Section 5.2, and compare with Trix without noise injection and Flock. To minimize the variance induced by injected noise and to ensure a fair comparison, we report ensembled prediction results with 16 samples for both Trix + noise and Flock.
(a) Zero-shot performance for entity prediction.
@@
(b) Zero-shot performance for relation prediction.

### v2

@@ B.2 Invariance
Similarly, let $\eta=v_{0}\xrightarrow{r_{1}}\dots\xrightarrow{r_{\ell}}v_{\ell}$ be a walk of length $\ell$ in $G$. The view of $\eta$ with $\mu$ is defined as:
$\mu\left(v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{\ell}}v_{\ell}\right)=\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\dots\xrightarrow{\phi(r_{\ell})}\pi(v_{\ell})$
@@
Following the definition of invariance in probability, provided in Section˜3, we extend all the definitions above to the stochastic case, replacing equality (==) with equality in distribution ($\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}$).
We can now prove the main propositions stated in Section˜4.2. Let’s begin with the more general:
@@
We can use the conclusion from Proposition 4.2 to prove the probabilistic invariance of the architecture proposed in Section˜4. To be able to apply it, we first need to verify the invariance of all used components, which we formalize in the following lemmas.
Lemma B.6.
The choice of the first step $v_{0}\xrightarrow{r_{1}}v_{1}$ of the uniform random walk algorithm described in Section˜A.1 is invariant.
@@
Let $G=(V,E,R)$ be a graph and let $H\simeq G$ be an isomorphic graph, with the isomorphism $\mu=(\pi,\phi)$ taking $G$ to $H$. Consider a link prediction query $q=(h,r,?)$ over $G$, and its identification $q^{\prime}=\mu(q)=(\pi(h),\phi(r),?)$. The goal is to show that when using $\eta$ described in Section˜A.1 for $(G,q)$ and $(H,q^{\prime})$, the first steps:
@@
Let $G=(V,E,R)$ be a knowledge graph, and let $\ell$ be the length of random walks. Let $H$ be a KG isomorphic to $G$ via the isomorphism $\mu=(\pi,\phi)$. We aim to show that:
$\mu(\eta(G,\ell))=\pi(V_{0})\xrightarrow{\phi(R_{1})}\pi(V_{1})\xrightarrow{\phi(R_{2})}\dots\xrightarrow{\phi(R_{\ell})}\pi(V_{\ell})\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}U_{0}\xrightarrow{S_{1}}U_{1}\xrightarrow{S_{2}}\dots\xrightarrow{S_{\ell}}U_{\ell}=\eta(H,\ell)$
Let $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{\ell}}v_{\ell}\in\mathcal{W}(G)$ be a walk of length $\ell$ over $G$. It suffices to show that the probability of sampling $\bar{\eta}$ from $G$ is identical to the probability of sampling $\mu(\bar{\eta})$ from $H$:
${\mathbb{P}}(\eta(G,\ell)=\bar{\eta})={\mathbb{P}}(\eta(H,\ell)=\mu(\bar{\eta}))$
To see this, let us expand the definitions of ${\mathbb{P}}(\eta(G,\ell)=\bar{\eta})$:
${\mathbb{P}}(\eta(G,\ell)=\bar{\eta})=$ ${\mathbb{P}}(V_{0}=v_{0})$ (10)
@@
$\cdot\prod^{\ell-2}_{i=0}{\mathbb{P}}(V_{i+2}=v_{i+2}\mid V_{i+1}=v_{i+1},V_{i}=v_{i})$
$\cdot\prod^{\ell-1}_{i=0}{\mathbb{P}}(R_{i+1}=r_{i+1}\mid V_{i+1}=v_{i+1},V_{i}=v_{i})$
and $P(\eta(H,\ell)=\mu(\bar{\eta}))$:
${\mathbb{P}}(\eta(H,\ell)=\mu(\bar{\eta}))=$ ${\mathbb{P}}(U_{0}=\pi(v_{0}))$ (11)
@@
$\cdot\prod^{\ell-2}_{i=0}{\mathbb{P}}(U_{i+2}=\pi(v_{i+2})\mid U_{i+1}=\pi(v_{i+1}),U_{i}=\pi(v_{i}))$
$\cdot\prod^{\ell-1}_{i=0}{\mathbb{P}}(S_{i+1}=\phi(r_{i+1})\mid U_{i+1}=\pi(v_{i+1}),U_{i}=\pi(v_{i}))$
@@
for all $i\in\{0,1,\dots,\ell-2\},j\in\{1,\dots,\ell-1\}$. Moreover, by the assumption that the first step $V_{0}\xrightarrow{R_{1}}V_{1}$ is invariant, we have:
@@
Substituting these equalities into (13) and multiplying both sides by the equalities from (12) for all choices of $i\in\{0,1,\dots,\ell-2\},j\in\{1,\dots,\ell-1\}$, we get precisely the equality of the right sides of Equation˜10 and Equation˜11. Hence,
${\mathbb{P}}(\eta(G,\ell)=\bar{\eta})={\mathbb{P}}(\eta(H,\ell)=\mu(\bar{\eta}))$
and we can conclude that $\mu(\eta(G,\ell))\stackrel{{\scriptstyle\mathclap{\text{{\mbox{$d$}}}}}}{{=}}\eta(H,\ell)$, and the algorithm $\eta$ is invariant in probability. ∎
Corollary B.8.
The random walk algorithm presented in Section˜A.1 is invariant in probability.
Lemma B.9.
The recording protocol $w$, as described in Section˜A.2, is invariant, provided that the embedding functions $\mathbf{v}$ and $\mathbf{r}$ are invariant.
@@
Let $G=(V,E,R)$ and $H=(V^{\prime},E^{\prime},R^{\prime})$ be isomorphic knowledge graphs with the isomorphism $\mu=(\pi,\phi)$ taking $G$ to $H$. Let $q=(h_{q},r_{q},?)$ be a link prediction query over $G$, and $\mu(q)=(\pi(h_{q}),\phi(r_{q}),?)$ be its identification in $H$. Let $\bar{\eta}=v_{0}\xrightarrow{r_{1}}v_{1}\xrightarrow{r_{2}}\dots\xrightarrow{r_{\ell}}v_{\ell}\in\mathcal{W}(G)$ be a walk over $G$, and $\bar{\eta}^{\prime}=\mu(\bar{\eta})=\pi(v_{0})\xrightarrow{\phi(r_{1})}\pi(v_{1})\xrightarrow{\phi(r_{2})}\dots\xrightarrow{\phi(r_{\ell})}\pi(v_{\ell})$ be the analogous walk over $H$. To prove that the recording protocol $w$ outlined in Section˜A.2 is invariant, it suffices to show that the encoding of each step:
@@
Lemma B.10.
The consensus protocol $c$, as described in Section˜A.4, is invariant.
@@
Flock with components as described in Section˜4 is invariant in probability.
@@
The result follows naturally from aggregating the results of Corollary B.8 and Lemmas B.9 and B.10, followed by applying Proposition 4.2. ∎
Appendix C Details of the Petals benchmark
Motivated by this limitation, we introduce the Petals benchmark. Petals comprises 220 graphs, each paired with a link prediction query $(h,r,?)$ and a target set $\{t_{1},t_{2}\}$. While $t_{1}$ and $t_{2}$ are non-isomorphic, KGFMs enforcing relational invariance are unable to distinguish them, producing identical predictions. We empirically validate this property by evaluating the classification accuracy of marking $t_{1}$ as true and $t_{2}$ as false, reported in Table˜1.
C.1 Structure of the studied KGs
Knowledge graphs in Petals follow a flower-like structure, parametrized by the number $c$ of ‘petals’, their length $l$ and the length $t$ of the ‘stem’ (see Figure˜5 for visualization).
Edges. The nodes of the stem are connected in a consecutive manner by the same relation type $r_{0}$. Precisely, for each $i\in{1,\cdots,t}$, there exists an edge $(b_{i-1},r_{0},b_{i})$. Each petal $A^{(i)}$ is associated with two edge types $r^{(i)}_{1},r^{(i)}_{2}$, and is connected to the central node $b_{0}$ with links $\left(b_{0},r^{(i)}_{1},a_{1}^{(i)}\right)$ and $\left(b_{0},r^{(i)}_{2},a_{2}^{(i)}\right)$. The rest of the petal is connected with edges of type $r_{1}^{(i)}$ only, going from $a^{(i)}_{2j-1}$ to $a^{(i)}_{2j+1}$, and from $a^{(i)}_{2j}$ to $a^{(i)}_{2j+2}$. Finally, there are also edges linking $a^{(i)}_{2\ell-1}$ and $a^{(i)}_{2l}$ to $a^{(i)}_{2l+1}$. Therefore, the full set of edges can be characterized as:
$\cup\left(\bigcup_{i=1}^{c}\bigcup_{j=1}^{\ell-1}\left\{\!\left(a^{(i)}_{2j-1},r^{(i)}_{1},a^{(i)}_{2j+1}\right),\!\left(a^{(i)}_{2j},r^{(i)}_{1},a^{(i)}_{2j+2}\right)\right\}\right)$
$\cup\left(\bigcup_{i=1}^{c}\left\{\!\left(a^{(i)}_{2\ell-1},r^{(i)}_{1},a^{(i)}_{2l+1}\right),\!\left(a^{(i)}_{2l},r^{(i)}_{1},a^{(i)}_{2l+1}\right)\right\}\right)$
We select each of the types $r^{(i)}_{1}$ and $r^{(i)}_{2}$ from the set of considered relations $R=\{r_{1},\dots,r_{|R|}\}$ so that any relation-invariant model will equate all petals (i.e. so that for each pair of petals, there is an automorphism taking one to another). For instance, Figure˜5 displays a cyclic pattern, in which $r^{(i)}_{2}=r^{(i+1)}_{1}$. Such symmetry causes all petals to be isomorphic, and leads to the inability of KGFMs to distinguish between the relations inside them.
Link prediction instances. Although the petals are isomorphic to each other, given the asymmetry of edge types from $b_{0}$ to $a^{(i)}_{1}$ and $a^{(i)}_{2}$, the nodes within a single petal generally can be distinguished. Therefore, for each graph with the structure as described above, we randomly sample one of the stem nodes $b_{s}$, and ask the link prediction query $(b_{s},r_{0},?)$. For the target nodes, we randomly select petal index $i$ and distance $j$ from the central node $b_{0}$, and consider the predictions for $a^{(i)}_{2j-1}$ and $a^{(i)}_{2j}$. For example, Figure˜5 shows the case when $b_{s}=b_{0}$, $i=1$ and $j=1$, where the query is $(b_{0},r_{0},?)$ and we are interested in the scores for $a^{(1)}_{1}$ and $a^{(1)}_{2}$.
C.2 Parameters and generation
@@ Appendix D Computational Complexity
Recall that $I$ is the iterations in each forward pass of Flock; $n$ is the base walk count; $\ell$ is the walk length; $L$ is the number of linear sequence-model layers (such as GRU); and $d$ is the hidden dimension for the sequence processor. Note that in practice, we perform $P$ forward passes and ensemble their outputs to reduce variance. For a single pass (P=1P{=}1), walk sampling and recording cost $O(n\ell)$, while the sequence processor with $L$ layers of hidden dimension $d$ costs $O(n\ell Ld^{2})$. The consensus protocol costs $O(n\ell d)$. In total, the time complexity is $O\big(PIn\ell Ld^{2}\big)$, which scales linearly with the number of (base) walks $n$, the length of walks $\ell$, and the number of ensembled predictions $P$. We empirically verified this in Appendix˜E.
@@ Appendix E Scalability analysis
Table 7: Inference scalability on a single NVIDIA RTX A6000 (48 GB) with batch size =8=8. Left columns specify base walks $n$ and ensembled passes $P$. Dashes indicate not applicable.
@@
To investigate the scalability of the proposed method Flock, we report the training and inference time per batch and peak GPU memory for Ultra, Trix, and variants of Flock on a single RTX A6000 (48 GB) in Tables˜6 and 7.
@@
During training, we fix Flock to $n=16$ base walks and with an ensemble size of $P=1$, which yields higher cost than Ultra/Trix but remains feasible on a single GPU. In addition, unlike Ultra/Trix, Flock does not rely on GNN message passing where highly optimized fused sparse kernels (e.g., RSPMM kernel developed in Zhu et al. (2021)) accelerate computation; instead, Flock’s runtime is dominated by walk sampling and sequence encoding, making time per batch the main bottleneck. As a result, pretraining typically takes about three days. One avenue for future work is to develop similarly highly optimized kernels for random-walk sampling.
@@
Additionally, we report the inference results in Table˜7, where we vary the number of walks $n$ and ensembled passes $P$. We observe near-linear growth of latency and VRAM with $n$, reflecting the dominant costs of walk sampling and sequence processing. Note that during inference, ensembled predictions are parallelizable, meaning that with sufficient GPU memory, these $P$ stochastic passes can be executed concurrently, so the latency grows sublinearly in $P$, while peak VRAM scales roughly linearly with $P$. In practice, reducing $n$ (walks) or $P$ (ensembled passes) lowers both memory and latency, while larger $n/P$ settings trade extra cost for better coverage and stability on harder KGs.
@@ Appendix F Noise injection over existing KGFMs
Table 8: Noise injection over the best performing KGFM baseline Trix.
(a) Zero-shot entity prediction.
@@
(b) Zero-shot relation prediction.
@@
(c) Accuracy on Petals.
Accuracy
Trix 50%
+ noise 52%
Flock 100%
Setup.
Since noise injection is a possible way to build a probabilistic equivariant KGFM in a different way from our approach (Gao et al., 2023), it is natural to ask how such KGFMs would perform compared to Flock. To answer this question, we apply noise injection over the best performing KGFMs baselines Trix. Specifically, in each forward pass, we add element-wise noise sampled from a uniform distribution $\epsilon\sim{\mathcal{U}}[-0.5,0]$ to all relation and entity embeddings after the initialization stage. Note that the addition of noise technically breaks deterministic node-relation equivariance, but the resulting model (Trix + noise) still respects probabilistic node-relation equivariance. We then pretrain Trix using the same experimental setup shown in Section˜5.2, and compare with Trix without noise injection and Flock. To minimize the variance induced by injected noise and to ensure a fair comparison, we report ensembled prediction results with 16 samples for both Trix + noise and Flock. This is a strong baseline implementing the ideas of prior work on noise injection and test-time ensembling for message passing networks on KGs (Lee et al., 2023; Gao et al., 2023).